the area of a circle is 9 pi square feet. what is the radius of the circle?

Answer by jim_thompson5910(35256)

A = pi*r^2

9pi = pi*r^2

9pi/pi = pi*r^2/pi

9 = r^2

r^2 = 9

r = sqrt(9)

r = 3


The radius is 3 feet
--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------